# disposition

### v2

@@ Abstract
It is well-known that scalarization techniques (e.g., in the sense of Gerstewitz; Kasimbeyli; Pascoletti and Serafini; Zaffaroni) are useful for generating (weakly, properly) efficient solutions of vector optimization problems. One recognized approach is the conic scalarization method in vector optimization in real normed spaces proposed by Kasimbeyli (2010, SIAM J Optim 20), which is based on augmented dual cones and Bishop-Phelps type (norm-linear) scalarizing functions. In this paper, we present new results on cone separation in real topological-linear spaces by using Bishop-Phelps type separating cones / separating seminorm-linear functions. Moreover, we study some extensions of known scalarization results in vector optimization (in the sense of Eichfelder; Gerstewitz; Jahn; Kasimbeyli; Pascoletti and Serafini). On this basis, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which can be seen as an extension of Kasimbeyli’s conic scalarization method in real normed spaces. Within this framework, we derive new Bishop-Phelps type scalarization results for the concepts of weak efficiency and different types of proper efficiency.
Abstract
Keywords: Vector optimization; (proper, weak) efficiency; Bishop-Phelps cone; augmented dual cone; Bishop-Phelps type scalarization; conic scalarization
Mathematics subject classifications (MSC 2010): 90C48, 90C29, 06F20, 52A05
@@ 1 Introduction
The field of vector optimization in real topological-linear spaces is concerned with the minimization of a vector-valued function $f:X\to Y$ acting between real topological-linear spaces $X$ and $Y$ with respect to a convex cone $C\subseteq Y$ over a nonempty feasible set $\Omega\subseteq X$. A review of the existing literature reveals that many papers focus on the (finite dimensional) normed setting (i.e., $Y$ is a normed space), see e.g. [16], [8, 9], [12, 13], [17], [32], [45, 46], [53], [58], [57, Ch. 7 and Sec. 15.2], and references therein. In particular, the analysis includes real normed spaces $Y$ where the convex cone $C$ is of Bishop-Phelps type (named after Bishop and Phelps [2], see also [16], [15], [28, 29], [36, 37], [35],[41, 43], [45, 46]). The importance of the class of Bishop-Phelps cones can be emphasized by the fact that a nontrivial, closed, convex cone in a real normed space is representable as a Bishop-Phelps cone if and only if it is well-based (see Example 2.4 for details). We are primarily interested on the more general setting of a real (possibly infinite dimensional and not normed) topological-linear space $Y$ motivated by optimization problems under uncertainty and problems from mathematical finance. Especially, robust counterparts to optimization problems under uncertainty can be formulated as vector optimization problems with an infinite dimensional topological-linear image space (see Göpfert et al. [23, Ch. 6] and references therein). Furthermore, in arbitrage theory (see the class of financial market models in Marohn and Tammer [51, 52], and Riedel [56]), problems in the setting of linear spaces are considered. Some further works in vector optimization that focus on this setting are [5], [20, 21], [22], [23], [30], [34], [25], [42, 43], [57].
@@
The main focus in our paper is on the computation of efficient / weakly efficient / properly efficient solutions of the given vector optimization problem via scalarization (based on Bishop-Phelps type cones), where the proper efficiency concepts considered in this paper cover certain known concepts of proper efficiency from the literature (e.g., in the sense of Benson [1]; Borwein [3, 4]; Henig [38]; Hurwicz [40]). It is well established that scalarization methods are useful for solving vector optimization problems. A notable feature of such methods is the ability to compute weakly efficient solutions by solving scalar optimization problems, which are often parameter-dependent, instead of the original vector optimization problem. Moreover, using scalarization, one can derive necessary and sufficient optimality conditions for the initial vector optimization problem, using generalized differentiation concepts for the new, scalarized objective function (see, e.g., [24], [32], [10], [33], [8], [34], [9]). The application of the scalarization methods for Henig dilating cones (enlarging convex cones) of the convex cone $C$ yields solutions that are properly efficient in the sense of Henig [38]. The relationships of the Henig proper efficiency concept to other known proper efficiency concepts have been well studied in the literature (see, e.g., [5, Sec. 2.4.3], [25, Ch. 3], [26], [49, Sec. 2.4] and references therein).
We will now recall some well-known scalarization methods in vector optimization. Of course, the linear scalarization method in vector optimization (respectively, the weighted sum method in multiobjective optimization) is the best known technique for generating weakly / properly efficient solutions of vector (multiobjective) optimization problems. It is known that this method only works well under cone-convexity type assumptions (see, e.g., [27]). For nonconvex vector optimization problems, there are some other scalarization methods frequently used in the literature that we recall below. In the framework of real (possibly infinite dimensional) normed spaces, the conic scalarization method by Kasimbeyli [45, 46] (see also [16]), the Drummond–Svaiter scalarization method [7] (see also [6]) and the Zaffaroni scalarization method [58] (see also [14, Sec. 5.1]) based on the Hiriart-Urruty function [39] (also known as oriented/signed distance function) are known to be useful, where the scalarizing function of the first two methods involves the underlying norm of the space $Y$. Considering the more general framework of real topological-linear spaces, only a few methods (besides the linear scalarization method) are known to be valid, such as the Gerstewitz scalarization method [20, 21] (see also [22], [49], [57], [23]) and the Pascoletti–Serafini scalarization method [53] (see also [12, 13], [57, Sec. 6.2.2]). Certain well-known scalarization methods in vector and multiobjective optimization (e.g. linear scalarization method / weighted sum method, $\varepsilon$-constraint method, hybrid method, Benson method, weighted Chebyshev norm method, and other methods; see [57, Sec. 7.3]) are special cases of these general scalarization methods. Some further comparison studies of scalarization methods in vector optimization are given in [6], [48], [12, Ch. 2], while generalizations of scalarization methods to vector and set optimization involving variable domination structures are also studied in the recent literature (see, e.g., [35], [15], [14], [43], [57, Ch. 10 and 11] and references therein).
@@
As mentioned above, one recognized approach in vector optimization in a real normed space $(Y,||\cdot||)$ is the conic scalarization method in vector optimization proposed by Kasimbeyli [45], which is based on the following parameterized scalar optimization problem:
$\begin{cases}x^{*}(f(x)-a)+\alpha||f(x)-a||\to\min\\ x\in\Omega\end{cases}$ (1.1)
@@
with a pair $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ (respectively, $(x^{*},\alpha)$ belongs to the so-called augmented dual cone of $C$; see Section 2.3 for details) and a given reference point $a\in Y$. Several relationships between (weakly, properly) efficient solutions of the original vector problem and solutions of scalar problems of type (1.1) are known (see [16], [17], [31], [45, 46]). In this paper, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which can be seen as an extension of Kasimbeyli’s scalarization method in real normed spaces. By taking a look on (1.1) one can see that Kasimbeyli’s method is based on Bishop-Phelps type norm-linear scalarizing functions (i.e., functions that are a nonnegative weighted sum of the norm |||||$\cdot$||| | ⋅ | | with a linear continuous functional $x^{*}$, see also Zaffaroni [59]), where the involving norm |||||$\cdot$||| | ⋅ | | is exactly the underlying norm in the real normed space $Y$. In our Bishop-Phelps type scalarization method, we consider Bishop-Phelps type seminorm-linear scalarizing functions (i.e., functions that are a nonnegative weighted sum of a seminorm $\psi$ with a linear continuous functional $x^{*}$), where the seminorm involved (namely $\psi$) must not have anything to do with the underlying topology of the real (locally convex) topological-linear space (e.g., the seminorm must not be part of a family of seminorms that generates this topology).
The rest of the paper is organized as follows:
In Section 2, after presenting some basics in real topological-linear spaces including properties of seminorms, we study properties of cones and their bases, focusing in particular on augmented dual cones, Bishop-Phelps type cones and zero level sets of Bishop-Phelps type (seminorm-linear) functions. We then introduce the solution concepts of vector optimization necessary for our work.
Section 3 is devoted to the study of scalarization methods in vector optimization which are based on cone monotone and cone representing scalarizing functions. Beside classical scalarization results (in the sense of Eichfelder [12, 13]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]), we establish some new scalarization results (namely Propositions 3.11, 3.14 and 3.17) for vector optimization problems in real topological-linear spaces, which extend the classical ones.
In Section 4, we present the key tool to prove our main Bishop-Phelps type scalarization results in vector optimization, namely the technique of nonlinear separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions. Besides presenting some classical nonlinear cone separation results, we also reflect on the separation conditions involved in both real topological-linear spaces and real normed spaces. Propositions 4.1 and 4.11 give some new insights into the relationships between the classical separation condition (the intersection of two cones is the singleton set $\{0\}$) and the separation conditions used in the nonlinear separation approach based on Bishop-Phelps type cones / functions discussed in [16], [28, 29], [45, 46]. We are able to deduce some new variants of nonlinear cone separation theorems involving the classical separation condition (see Corollaries 4.16 and 4.17).
@@
The main Section 5 is dedicated to Bishop-Phelps type scalarization for vector optimization in real topological-linear spaces. The parameterized scalarization problems (in the sense of Kasimbeyli [45]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]) are formulated using Bishop-Phelps type cone representing (seminorm-linear) functions. In addition to the basic scalarization results for the original vector optimization problem (roughly speaking, solutions of the scalar problems are solutions of the vector problem), we also point out (for the concept of efficiency in Proposition 5.12 and for the concept of weak efficiency in Proposition 5.14) the relationships between the original vector problem (based on the convex cone $C$), an updated vector problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions (e.g., seminorm-linear scalarizing functions or Gerstewitz [20, 21] scalarizing functions). Then, using nonlinear cone separation results (from Section 4) and Bishop-Phelps type scalarization results (Propositions 5.12 and 5.14), we are able to present new scalarization theorems (Theorem 5.19 for weak efficiency, Theorems 5.25, 5.38 and 5.40 for proper efficiency) and their corollaries, which show that any (weakly, properly) efficient solution of the original vector problem is a solution of the above mentioned (parameterized) scalar problems for certain choices of parameters.
Finally, in Section 6 we present a conclusion and an outlook for further research.
@@ 2.1 Basics in topological-linear spaces
Throughout the paper, we assume that $Y$ is a real topological-linear space. As usual, $Y^{*}$ denotes the topological dual space of $Y$. In certain situations we will also consider a more specific setting where $(Y,||\cdot||)$ is a real normed space. Consistently, the set of nonnegative real numbers is denoted by $\mathbb{R}_{+}$, while $\mathbb{P}:=\mathbb{R}_{++}$ denotes the set of positive real numbers. Given a set $\Omega\subseteq Y$, we denote by ${\mathrm{c}onv}\,\Omega$ the convex hull of $\Omega$. Moreover, we denote by ${\mathrm{c}l}\,\Omega$, ${\mathrm{b}d}\,\Omega$, and ${\mathrm{i}nt}\,\Omega$ the closure, the boundary, and the interior of $\Omega$ with respect to the underlying topology in $Y$. The algebraic interior (the core) of $\Omega$ is defined by $\operatorname{cor}\,\Omega:=\{x\in\Omega\mid\forall\,v\in Y\;\exists\,\varepsilon>0:\;x+[0,\varepsilon]\cdot v\subseteq\Omega\}.$ It is known that $\operatorname{int}\,\Omega\subseteq\operatorname{cor}\,\Omega\subseteq\Omega\subseteq\operatorname{cl}\,\Omega=(\operatorname{int}\,\Omega)\cup\operatorname{bd}\,\Omega,$ and if $\operatorname{int}\,\Omega\neq\emptyset$ (i.e., $\Omega$ is solid) and convex, then $\operatorname{cor}\,\Omega=\operatorname{int}\,\Omega$.
Given a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$, we define the $\psi$-unit ball by $\operatorname{\mathbb{B}}_{\psi}:=\{y\in Y\mid\psi(y)\leq 1\}$ and the $\psi$-unit sphere by $\operatorname{\mathbb{S}}_{\psi}:=\{y\in Y\mid\psi(y)=1\}.$
Remark 2.1
Suppose that $Y$ is a real topological-linear space and $\psi:Y\to\mathbb{R}$ is a seminorm. The following assertions hold:
  • $1^{\circ}$.
    $\psi$ is upper semicontinuous $\iff$ the set $\{y\in Y\mid\psi(y)<1\}$ is open $\iff$ $\psi$ is continuous at the origin $\iff$ $\psi$ is continuous.
  • $2^{\circ}$.
    $\psi$ is lower semicontinuous $\iff$ $\operatorname{\mathbb{B}}_{\psi}$ is closed.
@@
Remark that ${\mathrm{k}er}\,\psi=\{0\}$ if and only if $\psi$ is a norm, and the singleton set $\{0\}$ is closed if and only if the topological-linear space $Y$ is Hausdorff. Moreover, it is known (see, e.g., Gelfand [18, 19], Eberlein [11]) that if $Y$ is a Banach space and $\psi:Y\to\mathbb{R}$ is a lower semicontinuous seminorm (alternatively, the seminorm $\psi$ has the Baire property), then $\psi$ is continuous.
2.2 Cones and their bases
@@
A set $K\subseteq Y$ is called a cone if $0\in K=\operatorname{\mathbb{R}}_{+}\cdot K$. A cone $K\subseteq Y$ is said to be convex if $K+K=K$ (or equivalently, $K$ is a convex set); proper or nontrivial if $\{0\}\neq K\neq Y$; pointed if $\ell(K):=K\cap(-K)=\{0\}$. Given a convex cone $K\subseteq Y$ in the real topological-linear space $Y$, the set $\ell(K)$ is known as the lineality space of $K$. In this case, $K$ is a linear subspace of $Y$ if and only if $K=\ell(K)$, and moreover, $\operatorname{int}\,K=K+\operatorname{int}\,K$ and $K\setminus\ell(K)=K+(K\setminus\ell(K))$. If $K\subseteq Y$ is a cone, then $K\neq Y\iff 0\notin\operatorname{cor}\,K\iff 0\notin\operatorname{int}\,K,$ and if further $K$ is convex, then (see also [27, Lem. 2.10]) $K\neq Y\iff\ell(K)\cap\operatorname{int}\,K=\emptyset.$
@@
$K^{+}=(\operatorname{conv}\,K)^{+}=(\operatorname{cl}(\operatorname{conv}\,K))^{+}\text{ and }(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}=(\operatorname{conv}\,K)^{\#},$ (2.1)
but the inclusion $(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}$ can be strict (see Göpfert et al. [23, p. 55]).
In the following definition, we recall a general base concept for cones in real topological-linear spaces (cf. Göpfert et al. [23, Def. 2.1.42]).
Definition 2.2
@@
Remark 2.3
Taking into account (2.1), for any nontrivial cone $K\subseteq Y$ in a real normed space $Y$, it is well-known (see [23, Prop. 2.2.23 and 2.2.32]) that
$\operatorname{conv}\,K\text{ has a convex base}\iff K^{\#}\neq\emptyset,$ (2.2)

### v3

@@ Abstract
It is well-known that scalarization techniques (e.g., in the sense of Gerstewitz; Kasimbeyli; Pascoletti and Serafini; Zaffaroni) are useful for generating (weakly, properly) efficient solutions of vector optimization problems. One recognized approach is the conic scalarization method in vector optimization in real normed spaces proposed by Kasimbeyli (2010, SIAM J Optim 20), which is based on augmented dual cones and Bishop-Phelps type (norm-linear) scalarizing functions. In this paper, we present new results on cone separation in real topological-linear spaces by using Bishop-Phelps type separating cones / separating seminorm-linear functions. Moreover, we study some extensions of known scalarization results in vector optimization (in the sense of Eichfelder; Gerstewitz; Jahn; Kasimbeyli; Pascoletti and Serafini). On this basis, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which is based on Bishop-Phelps type cone representing and cone monotone scalarizing functions (e.g., Gerstewitz scalarizing functions or seminorm-linear scalarizing functions). Thus, our method also extends Kasimbeyli’s conic scalarization method from real normed spaces to real topological-linear spaces. Within this framework, we derive new Bishop-Phelps type scalarization results for the concepts of weak efficiency and different types of proper efficiency.
keywords:
Vector optimization, (proper, weak) efficiency, Bishop-Phelps cone, augmented dual cone, Bishop-Phelps type scalarization, conic scalarization
pacs:
MSC Classification]90C48, 90C29, 06F20, 52A05
@@ 1 Introduction
The field of vector optimization in real topological-linear spaces is concerned with the minimization of a vector-valued function $f:X\to Y$ acting between real topological-linear spaces $X$ and $Y$ with respect to a convex cone $C\subseteq Y$ over a nonempty feasible set $\Omega\subseteq X$. A review of the existing literature reveals that many papers focus on the (finite dimensional) normed setting and, in particular, the analysis includes real normed spaces $Y$ where the convex cone $C$ is of Bishop-Phelps type, see e.g. [1], [2, 3], [4, 5], [6], [7], [8, 9], [10], [11], [12, Ch. 7 and Sec. 15.2], [13], [14, 15], [16, 17], [18],[19, 20], and references therein. The importance of the class of Bishop-Phelps cones can be emphasized by the fact that a nontrivial, closed, convex cone in a real normed space is representable as a Bishop-Phelps cone if and only if it is well-based. We are primarily interested on the more general setting of a real (possibly infinite dimensional and not normed) topological-linear space $Y$ motivated by optimization problems under uncertainty and problems from mathematical finance. Especially, robust counterparts to optimization problems under uncertainty can be formulated as vector optimization problems with an infinite dimensional topological-linear image space (see Göpfert et al. [21, Ch. 6] and references therein). Furthermore, in arbitrage theory (see the class of financial market models in Marohn and Tammer [22, 23], and Riedel [24]), problems in the setting of linear spaces are considered. Some further works in vector optimization that focus on this setting are [25], [26, 27], [28], [21], [29], [30], [31], [32, 20], [12].
@@
The main focus in our paper is on the derivation of scalarization results for (weakly, properly) efficient solutions of the given vector optimization problem, where the proper efficiency concepts considered in this paper cover certain known concepts of proper efficiency from the literature (e.g., in the sense of Benson [33]; Borwein [34, 35]; Henig [36]; Hurwicz [37]). It is well established that scalarization methods are useful for solving vector optimization problems. A notable feature of such methods is the ability to compute weakly efficient solutions by solving scalar optimization problems, which are often parameter-dependent, instead of the original vector optimization problem. Moreover, using scalarization, one can derive necessary and sufficient optimality conditions for the initial vector optimization problem, using generalized differentiation concepts for the new, scalarized objective function (see, e.g., [38], [7], [39], [40], [2], [30], [3]). The application of the scalarization methods for Henig dilating cones (enlarging convex cones) of the convex cone $C$ yields solutions that are properly efficient in the sense of Henig [36]. The relationships of the Henig proper efficiency concept to other known proper efficiency concepts have been well studied in the literature (see, e.g., [25, Sec. 2.4.3], [31, Ch. 3], [41], [42, Sec. 2.4] and references therein).
@@
One recognized approach in vector optimization in a real normed space $(Y,||\cdot||)$ is the conic scalarization method in vector optimization proposed by Kasimbeyli [8], which is based on the following parameterized scalar optimization problem:
$\begin{cases}x^{*}(f(x)-a)+\alpha||f(x)-a||\to\min\\ x\in\Omega\end{cases}$ (1)
@@
with a pair $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ (respectively, $(x^{*},\alpha)$ belongs to the so-called augmented dual cone of $C$) and a given point $a\in Y$. By taking a look on (1) one can see that Kasimbeyli’s method is based on Bishop-Phelps type norm-linear scalarizing functions (i.e., functions that are a nonnegative weighted sum of the norm |||||$\cdot$||| | ⋅ | | with a linear continuous functional $x^{*}$, see also Zaffaroni [43]), where the involving norm |||||$\cdot$||| | ⋅ | | is exactly the underlying norm in the real normed space $Y$. Several relationships between (weakly, properly) efficient solutions of the original vector problem and solutions of scalar problems of type (1) are known (see [1], [6], [44], [8, 9]).
In this paper, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which is based on Bishop-Phelps type cone representing and cone monotone scalarizing functions (e.g., Gerstewitz scalarizing functions or Bishop-Phelps type seminorm-linear scalarizing functions, i.e., functions that are a nonnegative weighted sum of a seminorm with a linear continuous functional). The underlying scalarizations of the vector problem will be of the types (in the sense of Gerstewitz [26, 27]; Jahn [20]; Pascoletti and Serafini [10]):
$\begin{cases}\varphi(f(x)-a)\to\min\\ x\in\Omega\end{cases}$ ($P_{\varphi}^{a}$)
and
$\begin{cases}\lambda\to\min\\ \varphi(f(x)-a-\lambda k)\leq 0,\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$ ($P_{\varphi}^{a,k}$)
where $a\in Y$, $k\in Y\setminus\{0\}$, and $\varphi:Y\to\mathbb{R}$ is a Bishop-Phelps (type) cone representing function, i.e., the level set
$C_{\rm BP}:=-\{y\in Y\mid\varphi(y)\leq 0\}$
is a Bishop-Phelps (type) cone in $Y$. In Kasimbeyli’s conic scalarization method based on the scalar problem (1), the set
$C_{\rm BP}=-\{y\in Y\mid x^{*}(y)+\alpha||y||\leq 0\}$
with $(x^{*},\alpha)\in(Y^{*}\setminus\{0\})\times(\mathbb{R}_{+}\setminus\{0\})$ is indeed a Bishop-Phelps cone in the real normed space $Y$. Hence, our method can be seen as an extension of Kasimbeyli’s scalarization method from real normed spaces to real topological-linear spaces.
@@
The rest of the paper is organized as follows: In Section 2, after presenting some basics in real topological-linear spaces including properties of seminorms, we study properties of cones and their bases, focusing in particular on augmented dual cones, Bishop-Phelps type cones and zero level sets of Bishop-Phelps type (seminorm-linear) functions. We then introduce the solution concepts of vector optimization necessary for our work. Section 3 is devoted to the study of scalarization methods in vector optimization which are based on cone monotone and cone representing scalarizing functions. Beside classical scalarization results (in the sense of Eichfelder [4, 5]; Gerstewitz [26, 27]; Jahn [20]; Pascoletti and Serafini [10]), we establish some new scalarization results (namely Propositions 9, 10 and 11) for vector optimization problems in real topological-linear spaces, which extend the classical ones. The main Section 4 is dedicated to Bishop-Phelps type scalarization for vector optimization in real topological-linear spaces. The parameterized scalarization problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$) are formulated using Bishop-Phelps type cone representing and cone monotone scalarizing functions (e.g., Gerstewitz [26, 27] scalarizing functions or seminorm-linear scalarizing functions). In addition to the basic scalarization results for the original vector optimization problem (roughly speaking, solutions of the scalar problems are solutions of the vector problem), we also point out (for the concept of efficiency in Proposition 17 and for the concept of weak efficiency in Proposition 18) the relationships between the original vector problem (based on the convex cone $C$), an updated vector problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar problems involving Bishop-Phelps type cone representing and cone monotone scalarizing functions. In Section 4.2, we present the key tool to prove our main Bishop-Phelps type scalarization results in vector optimization, namely the technique of nonlinear separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions. Besides presenting some classical nonlinear cone separation results, we also reflect on the separation conditions involved in both real topological-linear spaces and real normed spaces. Propositions 19 and 23 give some new insights into the relationships between the classical separation condition (the intersection of two cones is the singleton set $\{0\}$) and the separation conditions used in the nonlinear separation approach based on Bishop-Phelps type cones / functions discussed in [1], [14, 15], [8, 9]. Then, using nonlinear cone separation results (from Section 4.2) and Bishop-Phelps type scalarization results (Propositions 17 and 18), we present new scalarization theorems (Theorem 29 for weak efficiency, Theorems 31, 36 and 37 for proper efficiency) and their corollaries, which show that any (weakly, properly) efficient solution of the original vector problem is a solution of the above mentioned (parameterized) scalar problems for certain choices of parameters. Finally, in Section 5 we present a conclusion and an outlook for further research.
@@ 2 Preliminaries
Throughout the paper, we assume that $Y$ is a real topological-linear space. As usual, $Y^{*}$ denotes the topological dual space of $Y$. In certain situations we will also consider a more specific setting where $(Y,||\cdot||)$ is a real normed space. Consistently, the set of nonnegative real numbers and the set of positive real numbers is denoted by $\mathbb{R}_{+}$ and $\mathbb{P}$, respectively. Given a set $\Omega\subseteq Y$, we denote by ${\rm conv}\,\Omega$ the convex hull of $\Omega$. Moreover, we denote by ${\rm cl}\,\Omega$, ${\rm bd}\,\Omega$, and ${\rm int}\,\Omega$ the closure, the boundary, and the interior of $\Omega$ with respect to the underlying topology in $Y$. The algebraic interior (the core) of $\Omega$ is defined by $\operatorname{cor}\,\Omega:=\{x\in\Omega\mid\forall\,v\in Y\;\exists\,\varepsilon>0:\;x+[0,\varepsilon]\cdot v\subseteq\Omega\}.$
@@
Given a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$ (i.e., $\psi$ is sublinear and symmetric), we define the $\psi$-unit ball by $\operatorname{\mathbb{B}}_{\psi}:=\{y\in Y\mid\psi(y)\leq 1\}$ and the $\psi$-unit sphere by $\operatorname{\mathbb{S}}_{\psi}:=\{y\in Y\mid\psi(y)=1\}.$ In what follows, we will suppose that $\psi$ is nontrivial (i.e., $\psi\not\equiv 0$). It is well-known that if $Y$ is a real topological-linear space and $\psi:Y\to\mathbb{R}$ is a seminorm, then $\psi$ is upper semicontinuous $\Longleftrightarrow$ $\{y\in Y\mid\psi(y)<1\}$ is open $\Longleftrightarrow$ $\psi$ is continuous at the origin $\Longleftrightarrow$ $\psi$ is continuous. Moreover, $\psi$ is lower semicontinuous $\Longleftrightarrow$ $\operatorname{\mathbb{B}}_{\psi}$ is closed. Remark also that $\{y\in Y\mid\psi(y)=0\}=\{0\}$Longleftrightarrow $\psi$ is a norm. The singleton set $\{0\}$ is closed $\Longleftrightarrow$ the topological-linear space $Y$ is Hausdorff. Moreover, it is known (see, e.g., Gelfand [45, 46], Eberlein [47]) that if $Y$ is a Banach space and $\psi:Y\to\mathbb{R}$ is a lower semicontinuous seminorm, then $\psi$ is continuous.
@@
A set $K\subseteq Y$ is called a cone if $0\in K=\operatorname{\mathbb{R}}_{+}\cdot K$. A cone $K\subseteq Y$ is said to be convex if $K+K=K$ (or equivalently, $K$ is a convex set); proper or nontrivial if $\{0\}\neq K\neq Y$; pointed if $\ell(K):=K\cap(-K)=\{0\}$. Given a convex cone $K\subseteq Y$ in the real topological-linear space $Y$, the set $\ell(K)$ is known as the lineality space of $K$. In this case, $K$ is a linear subspace of $Y$Longleftrightarrow $K=\ell(K)$, and moreover, $\operatorname{int}\,K=K+\operatorname{int}\,K$ and $K\setminus\ell(K)=K+(K\setminus\ell(K))$. If $K\subseteq Y$ is a cone, then $K\neq Y\iff 0\notin\operatorname{cor}\,K\iff 0\notin\operatorname{int}\,K,$ and if further $K$ is convex, then (see also [48, Lem. 2.10]) $K\neq Y\iff\ell(K)\cap\operatorname{int}\,K=\emptyset.$
@@
$K^{+}=(\operatorname{conv}\,K)^{+}=(\operatorname{cl}(\operatorname{conv}\,K))^{+}\text{ and }(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}=(\operatorname{conv}\,K)^{\#},$ (2)
but the inclusion $(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}$ can be strict (see Göpfert et al. [21, p. 55]).
In the following definition, we recall a general base concept for cones in real topological-linear spaces (cf. Göpfert et al. [21, Def. 2.1.42] for the case of a convex cone).
Definition 1 (Base).
@@
Remark 1.
Taking into account (2), for any nontrivial cone $K\subseteq Y$ in a real normed space $Y$, it is well-known (see [21, Prop. 2.2.23 and 2.2.32]) that
$\operatorname{conv}\,K\text{ has a convex base}\iff K^{\#}\neq\emptyset,$ (3)